OpenStudy (anonymous):

Help please ! Katherine uses her credit card to purchase a new television for $709.15. She can pay off up to $350 per month. The card has an annual rate of 25.7% compounded monthly. How much will she pay in interest? $23.90 $28.88 $8.05 $33.05

OpenStudy (anonymous):

25.7% is the annual interest and she can pay it off in less than a year so you need to find the monthly interest

OpenStudy (anonymous):

so 25.7/ 12 will give you the monthly interest

OpenStudy (anonymous):

see where i coming from ?

OpenStudy (anonymous):

O.o soo 0.0214166667

OpenStudy (anonymous):

yup thaz your monthly interest

OpenStudy (anonymous):

and you can kno how many months she will take by dividing the 709.15 by 350

OpenStudy (anonymous):

and there you go .,

OpenStudy (anonymous):

I got 2.0261428571

OpenStudy (anonymous):

yea ., well ., you kno that ,. its more than 2 times ., that she has to pay

OpenStudy (anonymous):

so ., she has to pay interest for one more month .,

OpenStudy (anonymous):

so its gonna be .,

OpenStudy (anonymous):

I don't really know . Im new to this course

OpenStudy (anonymous):

0.02 which is your monthly interest /100 * 709.15

OpenStudy (anonymous):

times 3

OpenStudy (anonymous):

becuz you paying for 3 months .,

OpenStudy (anonymous):

I got 0.4254

OpenStudy (anonymous):

25.7 divided by 12 is 2.14 whichi is your monthly interest

OpenStudy (anonymous):

so ., that 2.14 %/100% = 0.0214

OpenStudy (anonymous):

times 700.19 = 14.984

OpenStudy (anonymous):

times 3 gives me 44.952

OpenStudy (anonymous):

sry make i made mistake with the 700 valuue ., but all you have to do is change it to 709.15

OpenStudy (anonymous):

I dont get it .

OpenStudy (anonymous):

price of tv = 709.15 what she pays every month = 350 rate per year = 25.7% rate per month = 25.7 / 12 months = 2.14 how many months does she have to pay ? 709.15/ 350 = 2.02 that means she has to pay interest for 3 months becuz it has gone over 2 she has to pay for the .02 of the 2.02 so its three months interest so we have now 3 times 2.14/100 % times 709.15 ., see ., i broke it down

OpenStudy (anonymous):

45.52743 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

i dont understand how that answer and it's not in my answer choices o.O

OpenStudy (anonymous):

o crap ., um if you do the interest on the 350 ., you get the 8 . sum as your asnwer. ,sry

OpenStudy (anonymous):

c can be your answer.,

OpenStudy (anonymous):

but the way i looked at it ., they like didnt specify ., if you were paying interest on the 350 or the 709.15 .,

OpenStudy (anonymous):

please,., forgive me ., i was making horrible mistake .,

OpenStudy (anonymous):

ok ., um ., the 709.15 /350 = 2.026

OpenStudy (anonymous):

you use that ., and then ., say 2.026 /100 times 709.15 times 3 ., that should wrk

OpenStudy (anonymous):

interest is = prt/100

OpenStudy (anonymous):

um ., ok ., finally got it ., the problem was compound interest